OpenStudy (dmezzullo):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30. intercept = (-2, 0); y-intercept = (0, 6) intercept = (0, -6); y-intercept = (2, 0) x-intercept = (2, 0); y-intercept = (0, -6) intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (dmezzullo):

@faithelizabeth

OpenStudy (dmezzullo):

@darkb0nebeauty

OpenStudy (dmezzullo):

@mikala1

OpenStudy (anonymous):

do u really want it explained

OpenStudy (anonymous):

15x-5y = 30 -5y = 30-15x y = 3x-6 y = 3(0)-6 y = 0-6 y = -6, (0,6) (0) = 3x-6 0 = 3x-6 -3x = -6 x = 2, (2,0) Solution: x-intercept = (2,0); y-intercept = (0,-6)

OpenStudy (anonymous):

dang nvm thank you here we go

OpenStudy (dmezzullo):

TYVM

OpenStudy (anonymous):

ywnp

OpenStudy (anonymous):

i feel smart now

OpenStudy (anonymous):

you are smarter than lol